OpenStudy (anonymous):

x to the 3 fourths power divided by x to the 1 sixth power

OpenStudy (anonymous):

\(\Huge \frac{x^{\frac{3}{4}}}{x^{\frac{1}{6}}}\)

OpenStudy (anonymous):

That?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Okay, so the rule is: \(\Large \frac{x^a}{x^b} = x^{a-b}\)

OpenStudy (anonymous):

Okay so: \(\Huge \frac{x^{\frac{3}{4}}}{x^{\frac{1}{6}}} = x^{\frac{3}{4}-\frac{1}{6}}\)

OpenStudy (anonymous):

i think she left...

OpenStudy (anonymous):

Thank you :) Sorry, I'm trying to help someone else with their work, also.

OpenStudy (anonymous):

Do you understand? Can you take it from there?

OpenStudy (anonymous):

Yes. (:

OpenStudy (anonymous):

Lovely =D